When 50 j of heat was supplied to the system, it did 30 j of work. What is the change in internal energy of the system?
Answer #1

Apply first law of thermodynamics

Q - W = dE

Q = +50 since it is inlet

W = 30 since it is done by the system

Q - W = dE

dE = 50-30 = 20 J